Healthy Foods' balance sheet shows a total of $25 million long-term debt with a coupon rate of 8.50%. The yield to maturity on this debt is 8.00%, and the debt has a total current market value of $27 million. The company has 10 million shares of stock, and the stock has a book value per share of $5.00. The current stock price is $20.00 per share, and stockholders' required rate of return, rs, is 12.25%. The company recently decided that its target capital structure should have 35% debt, with the balance being common equity. The tax rate is 24%. What is the effect of the capital structure change on the WACC of the firm? Please use excel
Answer #1

Old WACC

D/E = MV debt/(share*price)

=27/(10*20)

=0.135

D/A = D/(E+D)
D/A = 0.135/(1+0.135)
=0.1189
Weight of equity = 1-D/A
Weight of equity = 1-0.1189
W(E)=0.8811
Weight of debt = D/A
Weight of debt = 0.1189
W(D)=0.1189
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 8*(1-0.24)
= 6.08
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=6.08*0.1189+12.25*0.8811
WACC =11.52%

New WACC

Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
12.25 = Unlevered cost of equity+0.135*(Unlevered cost of equity-8)*(1-0.24)
Unlevered cost of equity = 11.85
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 11.85+0.53846*(11.85-8)*(1-0.24)
Levered cost of equity = 13.43


new D/E = (D/A)/(1-(D/A))

=0.35/(1-0.35) = 0.53846

Weight of equity = 1-D/A
Weight of equity = 1-0.35
W(E)=0.65
Weight of debt = D/A
Weight of debt = 0.35
W(D)=0.35
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 8*(1-0.24)
= 6.08
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=6.08*0.35+13.43*0.65
WACC =10.86%